electricfield in disk

Grade:12

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear  satish,

Let us calculate the Electric Field at a point P above the center of a charged disk with radius of R and a uniform surface charge density of $ \sigma$ as shown in below figure.

 

\includegraphics[scale=.8]{EfieldDisk.eps}

Starting with the general formula for a surface charge

$\displaystyle {\bf E} = \frac{1}{4 \pi \epsilon_0} \int \frac{\sigma(r')({\bf r} - {\bf r'}) da'}{\left \vert{\bf r} - {\bf r'} \right \vert^3}$ (1)

 

choose a coordinate system. A disk clearly lends itself to cylindrical coordinates. As a refresher, the next figure shows the infinitesimal displacement, where we have the infinitesmal area $ da'$

cartesian coordinates:

$\displaystyle da' = dx'dy'$

cylindrical coordinates:

$\displaystyle da' = s'ds'd\phi'$

 

\includegraphics[scale=.5]{InfDis.eps}

The vectors to the source and field points that are needed for the integration in cylindrical coordinates

$\displaystyle {\bf r} = z \hat{z}$

$\displaystyle {\bf r'} = s' \hat{\phi}'$

therefore

$\displaystyle {\bf r} - {\bf r'} = z \hat{z} - s' \hat{\phi}'$

$\displaystyle \left \vert{\bf r} - {\bf r'} \right \vert = \sqrt{s'^2 + z^2} $

substituting these relationships into (1) gives us

$\displaystyle {\bf E} = \frac{\sigma}{4 \pi \epsilon_0} \int_0^{2\pi} \int_0^R ... ...}{\left ( s'^2 + z^2 \right )^{3/2}}\left ( z \hat{z} - s' \hat{\phi}' \right )$ (2)

 

As usual break up the integration into the $ z$ and $ \phi$ components

z component:

$\displaystyle {\bf E}_z = \frac{\sigma}{4 \pi \epsilon_0} \int_0^{2\pi} \int_0^R \frac{z \hat{z}s'ds'd\phi'}{\left ( s'^2 + z^2 \right )^{3/2}} $

Since $ \hat{z}$ is always in the same direction and has the same magnitude (unit vector), it is constant and can be brought out of the integration. Integrating the ds them

$\displaystyle {\bf E}_z = \frac{\sigma z \hat{z}}{4 \pi \epsilon_0} \int_0^{2\pi} d\phi' \int_0^R \frac{s'ds'}{\left ( s'^2 + z^2 \right )^{3/2}} $

using u substitution

$\displaystyle u = s'^2 + z^2$

$\displaystyle du = 2s' ds$

$\displaystyle ds = \frac{du}{2s}$

with the limits of integration becoming

$\displaystyle u(s'=0) = z^2$

$\displaystyle u(s'=R) = R^2 + z^2$

trasnforming the integral to

$\displaystyle {\bf E}_z = \frac{\sigma z \hat{z}}{4 \pi \epsilon_0} \int_0^{2\pi} d\phi' \int_{z^2}^{R^2 + z^2} \frac{u^{-3/2}du}{2} $

integrating

$\displaystyle {\bf E}_z = \frac{\sigma z \hat{z}}{4 \pi \epsilon_0} \int_0^{2\pi} d\phi' \left. \right \vert _{z^2}^{R^2 + z^2} -u^{-1/2}du$

evaluating the limits

$\displaystyle {\bf E}_z = \frac{\sigma z \hat{z}}{4 \pi \epsilon_0} \int_0^{2\pi} \left ( \frac{1}{z} - \frac{1}{\sqrt{R^2 + z^2}} \right ) d\phi' $

integrating again simply gives

$\displaystyle {\bf E}_z = \frac{\sigma z \hat{z}}{2 \epsilon_0} \left ( \frac{1}{z} - \frac{1}{\sqrt{R^2 + z^2}} \right ) $

$ {\bf\phi}$ component:

$\displaystyle {\bf E}_{\phi} = \frac{\sigma}{4 \pi \epsilon_0} \int_0^{2\pi} \int_0^R -\frac{s'^2\hat{\phi}'ds'd\phi'}{\left ( s'^2 + z^2 \right )^{3/2}} $

If you cannot simply see how the $ \phi$ component is zero through symmetry, then carry out the integration. The key thing to learn here, and why it is not good to just skip over the $ \phi$ component, is to realize that $ \hat{\phi}$ is not constant throughout the integration. Therefore, one cannot bring it out of the integration. What needs to be done is to substitute in for $ \hat{\phi}$ . An important result from cylindrical coordinates is the relation between its unit vectros and those of cartesian coordinates.

$\displaystyle \hat{s} = \cos \phi \hat{x} + \sin \phi \hat{y} $

$\displaystyle \hat{\phi} = -\sin \phi \hat{x} + \cos \phi \hat{y} $

$\displaystyle \hat{z} = \hat{z} $

Plugging in the $ \hat{\phi}'$ into our integral

$\displaystyle {\bf E}_{\phi} = \frac{\sigma}{4 \pi \epsilon_0} \int_0^{2\pi} \i... ...{x} + \cos \phi' \hat{y} \right )ds'd\phi'}{\left ( s'^2 + z^2 \right )^{3/2}} $

$ {\bf x}$ component:

To make our job easier, let us first integrate $ d\phi'$

$\displaystyle {\bf E}_{\phi}^x = \frac{\sigma \hat{x}}{4 \pi \epsilon_0} \int_0... ...ac{s'^2 ds'}{\left ( s'^2 + z^2 \right )^{3/2}} \int_0^{2\pi} \sin \phi' d\phi'$

Note how $ \hat{x}$ can be taken out of integral, so we get

$\displaystyle {\bf E}_{\phi}^x = \frac{\sigma \hat{x}}{4 \pi \epsilon_0} \int_0... ...{\left ( s'^2 + z^2 \right )^{3/2}} \left. \right \vert _0^{2\pi} - \cos \phi' $

Evaluating the limits, gives us the result we expected.

$\displaystyle {\bf E}_{\phi}^x = \frac{\sigma}{4 \pi \epsilon_0} \int_0^R \frac... ...\hat{x} ds'}{\left ( s'^2 + z^2 \right )^{3/2}} \left ( -1 - (-1) \right ) = 0 $

$ {\bf y}$ component:

$\displaystyle {\bf E}_{\phi}^y = \frac{\sigma \hat{y}}{4 \pi \epsilon_0} \int_0... ...c{s'^2 ds'}{\left ( s'^2 + z^2 \right )^{3/2}} \int_0^{2\pi} -\cos \phi' d\phi'$

integrating

$\displaystyle {\bf E}_{\phi}^y = \frac{\sigma \hat{y}}{4 \pi \epsilon_0} \int_0... ...}{\left ( s'^2 + z^2 \right )^{3/2}} \left. \right \vert _0^{2\pi} -\sin \phi' $

which once again yeilds a zero.

$\displaystyle {\bf E}_{\phi}^y = \frac{\sigma \hat{y}}{4 \pi \epsilon_0} \int_0^R \frac{s'^2 ds'}{\left ( s'^2 + z^2 \right )^{3/2}} \left ( 0 - 0 \right ) = 0 $

Since the x and y components are zero

$\displaystyle {\bf E}_{\phi} = 0$

Therefore, for a charged disk at a point above the center, we have

$\displaystyle {\bf E} = \frac{\sigma z \hat{z}}{2 \epsilon_0} \left ( \frac{1}{z} - \frac{1}{\sqrt{R^2 + z^2}} \right ) $

and rearranging

$\displaystyle {\bf E} = \frac{\sigma }{2 \epsilon_0} \left ( 1 - \frac{z}{\sqrt{R^2 + z^2}} \right ) \hat{z} $

 Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

Askiitians Expert

Ajit Singh Verma IITD